Sketch the region of integration, carefully with appropriate labels, and then perform the following integral: ∫(0 to pi/4)∫(0 to 4secO) r dr dO
1
Expert's answer
2020-11-09T13:25:07-0500

"\int _0^{\frac {\pi}{4}}d \theta \int_0^{4\sec \theta}rdr" "=\frac {1}{2}\int_0^{\frac{\pi}{4}}16\sec^2\theta d\theta" "=[\sec \theta =\frac{1}{\cos\theta}]" "=8\tan\theta|_0^{\frac {\pi}{4}}" "=8" The sketch is attached below.